OpenStudy (anonymous):

Sebuah pompa berbentuk silinder mempunyai diameter bagian dalam 2,59 cm dan panjang 50 cm. Saat berada dalam posisi horizontal, pompa ditekan dengan gaya 0,284 N di salah satu ujungnya untuk mendorong piston, sehingga pompa mengeluarkan udara dengan laju 30 m/s melalui lubang berdiameter 0,500 cm di ujung yang lain. Berapakah kecepatan piston saat itu? Anggap jejari piston sama dengan jejari bagian dalam pompa. Gunakan percepatan gravitasi 9,8 . caranya gimana kaka ??

OpenStudy (anonymous):

@tyzaQueen speak in english, please! I don't understand what you're talking about.

OpenStudy (anonymous):

Sikit sukar nak faham fizik dalam BI/BM kak. Boleh terjemah jadi English dulu? Pompa itu ape? :)

OpenStudy (anonymous):

haha hei @Shadowys , ur malaysian??

OpenStudy (anonymous):

Yup! Thivitaa didn't told you? :)

OpenStudy (anonymous):

knp u tak cakap dari awal..,?? yes she didn't

OpenStudy (anonymous):

@tyzaQueen Are you familiar with Bernoulli's principle? @gerryliyana Sorry bro lol Where are you from?

OpenStudy (anonymous):

hehe..., i'm from jakarta..,

OpenStudy (anonymous):

and u, where're u from?? selangor??

OpenStudy (anonymous):

Lol Ah. Across the seas. lol I'm from Johor. Is pompa a 'pump' in BI? I'm guessing her question is one of Bernoulli's principles

OpenStudy (anonymous):

yes..., it is

OpenStudy (anonymous):

this is my questions :) i am sorry for that :D A gas cylinder has an inner diameter of 2.59 cm and a length of 50 cm. When in the horizontal position, the pump is pressed with force 0.284 N at one end to push the piston, so remove the air pump at a rate of 30 m / s through a 0.500 cm diameter hole at the other end. What is the speed of the piston at the time? Assume the same faint faint piston inside the pump. Use the gravitational acceleration 9.8 m / s

OpenStudy (anonymous):

 Are you familiar with the equation of continuity? \(A_1 V_1 = A_2 V_2 \)

OpenStudy (anonymous):

If you are, then sub, \(A_1 = (2.59)^2 \pi , A_2=(0.5)^2 \pi, v_2 =30\) and voila!